# Genixer: Empowering Multimodal Large Language Model as a Powerful Data Generator

### v2

@@ 3.4 Automatic Data Generation and Filtering
Here, we introduce how we use Genixer and Genixer to generate data and then present two effective data filtering strategies for Genixer and Genixer, respectively. This is important and necessary for ensuring the generated data quality and thus enhancing the performance.
Data Generation and Filtering of Genixer on The General Tasks. The diversity of image resources and effective data filtering are two key steps to obtaining high-quality instruction tuning data. Here, we first describe the image sources used for the data generation with Genixer. To prosper the diversity of current instruction tuning data, we utilize the mixed dataset comprising 558K images from the LAION [43], CC3M [44], and SBU [39], as described in [29], and further adopt the 830K images from the original SBU datasets. These different image resources can help to generate more diverse instruction tuning data. We directly feed these 1.4M images to the trained Genixer and then designate the specific instruction = “This is the Common VQA task”, and finally we produce the 1.4M raw VQA triplets pending for the next filtering.
To assess the quality of general task data, we design a Fuyu-driven data filtering framework to automatically and effectively filter the samples generated by Genixer. As illustrated in Fig. 6, we design the text prompt as follows:
This framework is applicable to filter the four general tasks defined in Sec. 3.1. To assess the data quality of Common VQA and Adv VQA tasks, we substitute the variables and with the generated questions and answers, respectively. In the case of MC VQA evaluation, we replace the option letter (e.g., “B”) with the corresponding option content (e.g., “Jumping”) and then convert the format to match that of Common VQA for processing by Fuyu-8B [5], as shown in Fig. 6. For the MD type, we decompose multi-turn dialogue into individual single-turn VQA instances, processing these in the same manner as Common VQA.
To avoid the potential bias to output “Yes” of Fuyu-8B, instead of directly getting the words Yes or No as the final results, we additionally calculate the probability of predicting the “Yes”. To this end, we restrict the output to either Yes or No and calculate the probability via the output logits as follows:
@@
where is the sign of keeping or removing the current sample, and denotes the probability of the result “Yes” of -th visual-question-answer triplet. By setting , we filter the 1.4M raw VQA triplets to 915K instances. We name this VQA-like synthetic dataset Genixer-915K.
Data Generation and Filtering of Genixer on The Grounding Tasks. Similarly, our Genixer adopts the same image resources in general tasks for generating grounding-based instruction tuning data. After feeding the 1.4M image corpus to Genixer by set the data type as REC, we get 1.4M raw data.
@@ 4 Experiments
In Sec. 3.4, we use two trained generators Genixer and Genixer to generate Genixer-915K and Genixer-350K for general and grounding tasks, respectively. Here, we evaluate the quality of these data from several aspects, including statistical analysis, human evaluation, evaluation via training MLLMs, and ablation study.
@@ 4.1 Statistical Analysis
To evaluate the generation quality of Genixer, we conduct a comparative analysis using the VQAv2 [15] training set as a comparison. Fig. 7 (a) illustrates the distribution of question lengths of VQAv2 and Genixer-915K. Notably, our dataset exhibits a significant long tail, indicating a higher proportion of longer sentences compared to VQAv2. Additionally, the distribution of nouns and verbs, depicted in Fig. 7 (b) and (c), showcases the diverse vocabulary present in the generated questions. To assess data quality, we utilize Flickr30K [52] as the image source and employ Genixer to generate three representative data types, as detailed in Sec. 3.4. Results in Tab. 2 demonstrate that our generated data achieves an accuracy exceeding 80%, with the highest accuracy observed in the MC VQA category. Furthermore, Eq. (3) guarantees that our generated data exhibits a high probability (0.8) of being classified as “Yes” across all three data types. This is corroborated by the probability distribution depicted in Fig. 9, affirming the high quality of data produced by Genixer in generating diverse instruction tuning data.
The statistics of the generated dataset Genixer-350K by Genixer are presented in Fig. 8, showcasing metrics related to expression length, region width, and height. Tab. 4 offers a comparative analysis of our dataset, highlighting the larger collection of images and expression lengths in Genixer-350K compared to other grounding-based datasets.
@@ 4.2 Human Evaluation
We conduct the human evaluation to manually analyze the generated question type and corresponding correctness. We employ Genixer to generate QA samples without specific the data type. Since the image content can affect the generation types, we randomly chose 100 images from the COCO validation set as the held-in set and 100 images from Flickr30K as the held-out set. As shown in Tab. 4, we divide the questions into seven types, such as “Action” and “Color”. One can observe that “Object Type” and “Relative position” are the two most popular question types of both held-in and held-out datasets. Among held-in samples, “Action” question also has a notable proportion with 92% correctness. Different from the held-in set, 38 out of 100 samples belong to “Object Type” and 23 out of 100 “Relative Position” in the held-out set. In all of these types, the correctness of held-in dataset is slightly higher than held-out dataset.
@@ 4.3 Evaluation via Training MLLMs
Evaluation on General Tasks. Considering the data size (665K) is small that is used for training LLaVA1.5 and the data balance during the finetuning stage, we proceed to add the Genixer-915K dataset into the pretraining phase, adhering to the same training protocols utilized by LLaVA1.5. Furthermore, we enriched the fine-tuning stage with an additional 8K VQA-like data points, selected based on Fuyu-8B’s probability range of 0.5 to 0.7. This probability range was chosen based on findings that higher probabilities correlate with simpler, more straightforward VQA instances. By selecting data within this specific range, our model is able to learn more challenging VQA samples. From Tab. 5, one can observe consistent enhancements across 10 out of 12 benchmarks compared with vanilla LLaVA1.5. Indeed, on several tasks, our generated data can make significant improvements, e.g., 3.8% on VizWiz, 2.9% on SciencQA, and 37.7 scores on the MME benchmark.
@@
Qwen-VL [3] | Qwen-7B | 78.8 | 59.3 | 35.2 | 67.1 | 63.8 | – | – | 38.2 | 7.4 | 62.3 | – | – |
Qwen-VL-Chat [3] | Qwen-7B | 78.2 | 57.5 | 38.9 | 68.2 | 61.5 | – | 1487.5 | 60.6 | 56.7 | 65.4 | – | – |
LLaVA-1.5 | Vicuna-7B | 78.5 | 62.0 | 50.0 | 66.8 | 58.2 | 85.9 | 1465.0 | 64.3 | 58.3 | 66.2 | 65.4 | 31.1 |
LLaVA-1.5 | Vicuna-7B | 79.1 | 63.1 | 53.8 | 69.7 | 59.0 | 87.3 | 1502.7 | 65.3 | 59.4 | 66.6 | 64.0 | 30.1 |
@@
OFA-L* [48] | 79.96 | 83.67 | 76.39 | 68.29 | 76.00 | 61.75 | 67.57 | 67.58 | 72.65 |
OFASys [4] | 80.10 | – | – | – | – | – | – | – | – |
VisionLLM-H [50] | 86.70 | – | – | – | – | – | – | – | – |
@@
Evaluation on Grounding Tasks. Here, we adopt Shikra [9] as the model to evaluate the quality of Genixer-350K REC-like data produced by Genixer. Adhering to Shikra’s published code, we incorporate our synthetic data into the training phases, maintaining consistent training iterations to ensure a fair comparison. Tab. 6 shows the improvement on 7 out of 8 test datasets with a non-trivial average boost of 0.6%. These findings imply that our pipeline can be an alternative approach to generate grounding-based instruction tuning data, which is typically challenging for manually labeling and not satisfied for prompting GPT-4V, as shown in Fig. 2.
Performance on Genixer. As illustrated in Tab. 2 and Fig. 9, Genixer showcases a superior capability of generating high-quality data. It is natural for us to investigate the performance of Genixer. Accordingly, we evaluate Genixer and report its results in Tab. 7, which refers to the setting ZS. One can observe the minor declines in performance on the GQA, ScienceQA, and SEED, alongside a modest enhancement on VQAv2, VizWiz, and POPE. Such results are due to the exclusive training on generating instruction tuning data. Thus, for a fair comparison to investigate the capability of Genixer, we proceeded to retrain the Genixer with the mixture of the 665K finetuning dataset used in LLaVA1.5 and the datasets for training Genixer for one epoch following the same training protocols. The outcomes of this process are presented as MixT in Tab. 7, where we witnessed significant improvements across all six benchmarks.
Method |
Setting |
VQAv2 | GQA | VizWiz | SQA | POPE | SEED |
LLaVA1.5 |
78.5 | 62.0 | 50.0 | 66.8 | 85.9 | 66.2 |
Genixer |
ZS |
79.7 | 61.3 | 50.8 | 65.6 | 86.0 | 65.6 |
Genixer |
MixT |
80.2 | 63.1 | 54.1 | 67.1 | 87.5 | 67.1 |
@@ 5 Conclusion, Limitations, and Societal Impacts
In this paper, we introduce a novel automatic data generation pipeline called Genixer, designed to efficiently and affordably produce high-quality instruction tuning data by leveraging current MLLMs. We instantiate Genixer into two data-generative MLLMs, Genixer and Genixer, tailored to generate general and grounding instruction tuning data, respectively. To ensure the quality of the generated data, we propose two data filtering frameworks: Fuyu-driven and CLIP-driven. Finally, we contribute two instruction tuning datasets, Genixer-915K and Genixer-350K, targeting Common VQA and REC. Experimental results demonstrate that both generated datasets significantly enhance LLaVA1.5 and Shikra across various multimodal benchmarks, respectively.

### v6

@@ 3.4 Automatic Data Generation and Filtering
Here, we introduce the image sources used for data generation. Additionally, recognizing that there are no ready-made state-of-the-art models available for the VQA data filtering, we introduce two data filtering pipelines to remove the incorrect data samples, which is essential for subsequent model training.
Data Generation and Filtering of GenixerL on The Generic Tasks. We utilize the mixed dataset comprising 558K images from the LAION [60], CC3M [61], and SBU [54], as described in [40], and further adopt the 830K images from the original SBU datasets. We directly feed these 1.4M images to the trained GenixerL and then designate the specific instruction = “This is the Common VQA task”, and finally we produce the 1.4M raw data samples.
To assess the quality of generic task data, we design a Fuyu-driven data filtering framework to automatically filter the samples, which may include incorrect questions or incorrect answers. We design the text prompt as follows:
For filtering Common VQA and Adv VQA tasks, we substitute the variables and with the generated questions and answers, respectively. In the case of MC VQA evaluation, we replace the option letter (e.g., “B”) with the corresponding option content (e.g., “Jumping”) and then convert the format to match that of Common VQA for processing by Fuyu-8B [5], as shown in Fig. 5. For the MD task, we decompose multi-turn dialogue into individual single-turn instances for filtering.
Rather than prompt Fuyu-8B to directly output “Yes” or “No” as the filtering label, we calculate the probability of predicting the “Yes” as follows:
@@
where is the filter label representing keeping or removing the current sample. denotes the probability of the result “Yes” of -th candidates. By setting , we filter the 1.4M raw VQA triplets to 915K instances. We name this VQA-like synthetic dataset Genixer-915K.
Data Generation and Filtering of GenixerS on The Grounding Tasks. GenixerS adopts the same image resources in generic tasks for generating grounding-based instruction tuning data. After feeding the 1.4M image corpus to GenixerS by set the data type as REC, we get 1.4M raw data.
@@ 4 Experiments
In this section, we evaluate the quality of synthetic datasets from several aspects, including statistical analysis, human evaluation, evaluation via training MLLMs, ablation study, visualizations, and user study.
@@ 4.1 Statistical Analysis
To evaluate the generation quality of GenixerL, we conduct a comparative analysis using the VQAv2 [22] training set as a comparison. Fig. 6 (a) illustrates the distribution of question lengths of VQAv2 and Genixer-915K. Notably, our dataset exhibits a significant long tail, indicating a higher proportion of longer sentences compared to VQAv2. Additionally, the distribution of nouns and verbs, depicted in Fig. 6 (b) and (c), showcases the diverse vocabulary present in the generated questions. To assess data quality, we utilize Flickr30K [76] as the image source and employ GenixerL to generate three representative data types, as detailed in Sec. 3.4. Results in Tab. 2 demonstrate that our generated data achieves an accuracy exceeding 80%, with the highest accuracy observed in the MC VQA category. Furthermore, our generated data exhibits a high probability (0.8) of being classified as “Yes” across all three data types. This is corroborated by the probability distribution depicted in Fig. 8, affirming the high quality of data produced by GenixerL in generating diverse instruction tuning data.
The statistics of the synthetic dataset Genixer-350K are presented in Fig. 7, showcasing metrics related to expression length, region width, and height. Tab. 4 offers a comparative analysis of our dataset, highlighting the larger collection of images and expression lengths in Genixer-350K compared to other grounding-based datasets.
@@ 4.2 Human Evaluation
We conduct the human evaluation to manually analyze the generated question type and corresponding correctness. We employ GenixerL to generate QA samples without specific the data type. Since the image content can affect the generation types, we randomly chose 100 images from the COCO validation set as the held-in set and 100 images from Flickr30K as the held-out set. As shown in Tab. 4, we divide the questions into seven types, such as “Action” and “Color”. One can observe that “Object Type” and “Relative position” are the two most popular question types of both held-in and held-out datasets. Among held-in samples, “Action” question also has a notable proportion with 92% correctness. Different from the held-in set, 38 out of 100 samples belong to “Object Type” and 23 out of 100 “Relative Position” in the held-out set. In all of these types, the correctness of held-in dataset is slightly higher than held-out dataset.
| Data Type | Accuracy(%) | Average Prob. |
| Common VQA | 82.4 | 0.8186 |
| MC VQA | 87.8 | 0.8721 |
| MD | 82.5 | 0.8252 |
@@ 4.3 Evaluation via Training MLLMs
Evaluation on General Tasks.
Benchmarks. Following the baseline LLaVA1.5 [39], we evaluate the enhanced model on 12 multimodal datasets and benchmarks. We select five generic VQA datasets, including VQAv2 [22], GQA [26], VizWiz [23], ScienceQA [46], and TextVQA [62]. We test the multimodal benchmarks on MME [18], MMBench [41], SEED-Bench [35], LLaVA-Bench [40] and MM-Vet [77]. Additionally, we also test the model on hallucination benchmark POPE [37].
Main results. As pointed out in [17] and [39], the ratio of dataset mixture is crucial for the model training. We proceed to add the large-scale synthetic dataset Genixer-915K into the pretraining stage rather than the finetuning stage for a fair comparison, adhering to the same training protocols utilized by LLaVA1.5. For the finetuning stage, we meticulously curated 8K VQA-like data samples, selected based on Fuyu-8B’s probability range of 0.5 to 0.7 with the image resource from SBU [54]. The probability range was chosen based on findings that higher probabilities correlate with simpler, more straightforward VQA instances. By selecting data within this specific range, our model can learn more challenging VQA samples. From Tab. 5, one can observe consistent enhancements across 10 out of 12 benchmarks compared with vanilla LLaVA1.5. Indeed, on several tasks, our generated data can make significant improvements, e.g., 3.8% on VizWiz, 2.9% on SciencQA, and 37.7 scores on the MME benchmark. While the improvement in the SEED benchmark is not as obvious as we initially expected, this is primarily because the synthetic data belongs to the Common VQA type while many questions of SEED require outside knowledge to solve which is out of the scope of Genixer-915K. The performance on LLaVA-Bench and MM-Vet shows a slight decline, possibly due to their smaller scale and the utilization of GPT-4 as the evaluation metric. This could introduce additional uncertainty and potential biases into the results.
@@
| Qwen-VL [3] | Qwen-7B | 78.8∗ | 59.3∗ | 35.2 | 67.1 | 63.8 | – | – | 38.2 | 7.4 | 62.3 | – | – |
| Qwen-VL-Chat [3] | Qwen-7B | 78.2∗ | 57.5∗ | 38.9 | 68.2 | 61.5 | – | 1487.5 | 60.6 | 56.7 | 65.4 | – | – |
| LLaVA-1.5 | Vicuna-7B | 78.5∗ | 62.0∗ | 50.0 | 66.8 | 58.2 | 85.9 | 1465.0† | 64.3 | 58.3 | 66.2 | 65.4 | 31.1 |
| LLaVA-1.5 | Vicuna-7B | 79.1∗ | 63.1∗ | 53.8 | 69.7 | 59.0 | 87.3 | 1502.7 | 65.3 | 59.4 | 66.6 | 64.0 | 30.1 |
@@
| GPV-2 [29] | 51.59 | – | – | – | – | – | – | – | – |
| OFA-L [67] | 79.96 | 83.67 | 76.39 | 68.29 | 76.00 | 61.75 | 67.57 | 67.58 | 72.65 |
| Unified-IO [45] | 78.60 | – | – | – | – | – | – | – | – |
| OFASys [4] | – | 80.10 | – | – | – | – | – | – | – |
| VisionLLM-H [69] | – | 86.70 | – | – | – | – | – | – | – |
| UNITER [13] | 81.41 | 87.04 | 74.17 | 75.90 | 81.45 | 66.70 | 74.02 | 68.67 | 76.17 |
| VILLA [19] | 82.39 | 87.48 | 74.84 | 76.17 | 81.54 | 66.84 | 76.18 | 76.71 | 77.77 |
| UniTAB [73] | 86.32 | 88.84 | 80.61 | 78.70 | 83.22 | 69.48 | 79.96 | 79.97 | 80.89 |
| MDETR [28] | 86.75 | 89.58 | 81.41 | 79.52 | 84.09 | 70.62 | 81.64 | 80.89 | 81.81 |
@@
Evaluation on Grounding Tasks.
Benchmarks. Following the baseline model Shikra [10], we test the enhanced model on REC tasks. We use the test datasets RefCOCO [30], RefCOCO+ [30], RefCOCOg [49].
Main results. Here, we adopt Shikra [10] as the model to evaluate the quality of Genixer-350K produced by GenixerS. Adhering to Shikra’s published code, we incorporate our synthetic data into the training phases, maintaining consistent training iterations to ensure a fair comparison. Tab. 6 shows the improvement on 7 out of 8 test datasets with a non-trivial average boost of 0.6%. These findings imply that our pipeline can be an alternative approach to generate grounding-based instruction tuning data, which is typically challenging for manually labeling and not satisfied for prompting GPT-4V, as shown in Fig. 1.
Performance on Genixer. As illustrated in Tab. 2 and Fig. 8, GenixerL showcases a superior capability of generating high-quality data. It is natural for us to investigate the performance of GenixerL. Accordingly, we evaluate GenixerL and report its results in Tab. 7, which refers to the setting ZS. One can observe the minor declines in performance on the GQA, ScienceQA, and SEED, alongside a modest enhancement on VQAv2, VizWiz, and POPE. Such results are due to the exclusive training on generating instruction tuning data. Thus, for a fair comparison to investigate the capability of GenixerL, we proceeded to retrain the GenixerL with the mixture of the 665K finetuning dataset used in LLaVA1.5 and the datasets for training GenixerL for one epoch following the same training protocols. The outcomes of this process are presented as MixT in Tab. 7, where we witnessed significant improvements across all six benchmarks.
| Method | Setting | VQAv2 | GQA | VizWiz | SQA | POPE | SEED |
| LLaVA1.5 | - | 78.5 | 62.0 | 50.0 | 66.8 | 85.9 | 66.2 |
| GenixerL | ZS | 79.7+1.2 | 61.3-0.7 | 50.8+0.8 | 65.6-1.2 | 86.0+0.1 | 65.6-0.6 |
| GenixerL | MixT | 80.2+1.7 | 63.1+1.1 | 54.1+4.1 | 67.1+0.3 | 87.5+1.6 | 67.1+0.9 |
@@ 4.4 Ablation Study
4.5 Visualizations
We sample some examples of general tasks such as Common VQA, MC VQA, and MD in Fig. 10, 11, 12, respectively. The examples of grounding tasks are shown in Fig. 13 and 14. These visualizations demonstrate that our GenixerL and GenixerS have the exceptional capability of generating diverse instruction tuning data. The generated VQA triplets in Fig. 10 show that GenixerL is capable of generating diverse types of data such as color, Yes/No, and counting types. The generated samples of Fig. 11 show the ability to generate multi-choice VQA data at a high-quality level. As for generating MD data, we display four examples in Fig. 12. The results demonstrate the ability to generate long semantic sentences when answering the questions. Furthermore, we also compare the REC data generation ability with GPT-4V [53] in Fig. 15. It indicates that GPT-4V is suboptimal for generating grounding tasks such as REC.
4.6 User Study
We conducted a user preference study to evaluate the generation quality between Genixer and GPT-4V. For this study, we selected 12 samples, comprising 4 Common VQA samples, 3 MC VQA samples, and 5 REC samples. Fig. 9 summarizes the statistical analysis of 13 valid surveys. (1) The first seven columns (excluding the first one) reveal that GPT-4V was the primary preference among users, while between 20% to 40% of users preferred the data generated by Genixer. Notably, a significant number of users selected “Tie” indicating that the data quality we generated is comparable to that of GPT-4V. (2) In analyzing the last five columns, our generated REC samples were more favored than those of GPT-4V, as the evidence shown in Fig. 15.
@@ 5 Conclusion, Limitations, and Societal Impacts
In this paper, we introduce a novel automatic data generation pipeline called Genixer, designed to efficiently and affordably produce high-quality instruction tuning data by leveraging current MLLMs. We instantiate Genixer into two data-generative MLLMs, GenixerL and GenixerS, tailored to generate general and grounding instruction tuning data, respectively. To ensure the quality of the generated data, we propose two data filtering frameworks: Fuyu-driven and CLIP-driven. Finally, we contribute two instruction tuning datasets, Genixer-915K and Genixer-350K, targeting Common VQA and REC. Experimental results demonstrate that both generated datasets significantly enhance LLaVA1.5 and Shikra across various multimodal benchmarks, respectively.